OpenStudy (anonymous):

Two similar solids are shown below. What is the volume of the larger solid? 2048 ft3 1024 ft3 4096 ft3 6042 ft3

OpenStudy (anonymous):

@radar @RadEn

OpenStudy (raden):

your file cant be opened

OpenStudy (xmoses1):

Yeah RadEn is right

OpenStudy (anonymous):

there @Xmoses1 & RadEn

OpenStudy (radar):

The linear dimension doubled, so the volume will be 8 (or twice cubed) times as much.

OpenStudy (radar):

Do you follow that reasoning?

OpenStudy (anonymous):

kinda, I'm really confused

OpenStudy (radar):

The volume is the result of cubing the linear dimension of a cube. wait a minute in this case it was obtained by lXwXh (length times width times height) resulting in cubic units. Now I need to know the relationship between solid 1 and solid 2. Did all dimensions double? can you print the exact wording of the problem.

OpenStudy (anonymous):

Two similar solids are shown below. [PICTURE] What is the volume of the larger solid? ANSWERS: 2048 ft3 1024 ft3 4096 ft3 6042 ft3

OpenStudy (radar):

I have assumed that the dimensions doubled all the way around in my first answer, yes, the two solids would then remain similar.

OpenStudy (anonymous):

I did 512 * 2 = 1024 ? and 1024/2 and got 512 so is it 1024?

OpenStudy (xmoses1):

I would say that radar knows what he is doing, so i will stay out of it:)

OpenStudy (anonymous):

I believe its "1024 ft3"

OpenStudy (mathstudent55):

When you take a line and double its size, the new line is 2 times as long as the original line. When you take a square and double the length of the side, the new area is 2^2 = 4 times the original area. When you take a cube and you double the length of the side, the new volume is 2^3 = 8 times bigger than the original.